OpenStudy (anonymous):

In order to solve the literal equation A equals s squared for s, what would you do to the equation? Divide both sides by s. Subtract s squared from both sides. Take the square root of both sides. Square both sides.

OpenStudy (solomonzelman):

C

OpenStudy (solomonzelman):

I can teach you if you don't mind

OpenStudy (anonymous):

How do you get it?

OpenStudy (solomonzelman):

do you want to know why am I taking a square root of n=both sides?

OpenStudy (solomonzelman):

Look you have  they are asking to solve for s,

OpenStudy (solomonzelman):

what do you think they want you to do?

OpenStudy (solomonzelman):

any ideas?

OpenStudy (solomonzelman):

I'll give you another example similar to this.

OpenStudy (solomonzelman):

d=x^2-2 solve for x you want to set the equation equal to x. add 2 to both sides d=x^2-2 d+2=x^2-2+2 d-2=x^2 take square root of both sides √(d+2)=√(x^2) so the answer would be √(d+2)=x get it @Hannnnah? can you do you example, you don't even need to subtract anything there, just take a square root of both sides.

OpenStudy (solomonzelman):

d+2=x^2-2+2 d+2=x^2 ( wrote -) take square root of both sides

OpenStudy (solomonzelman):

A=s^2 (just recalling)

OpenStudy (solomonzelman):

solve for s. (set it to s=....

OpenStudy (anonymous):

I get it now thanks, and sorry I was on the phone with a teacher, but thanks for helping

OpenStudy (solomonzelman):

Hannnnah, it OK! I was replying to 3 other questions, so it's OK! As long as you get it.......

OpenStudy (anonymous):

Yea lol

OpenStudy (solomonzelman):

Good!